OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12

OpenStudy (lgbasallote):

-3-6/10 -x = -3/2 -9/10-x = -3/2 -9 = -3/2 (10-x) -9(2) = -3(10-x) -9(2)/-3 = 10-x 6 = 10 -x x = 10 -6 x =4